The rim of a basketball hoop is 10 feet above the ground. The free-throw line is 15 feet from the rim of the basket. If the eyes of a basketball player are 6 feet above the ground, what is the angle of elevation of the player's line of sight when shooting a free throw to the rim of the basket?

please show all work so i can understand better

Answer by stanbon(57342)

The rim of a basketball hoop is 10 feet above the ground. The free-throw line is 15 feet from the rim of the basket. If the eyes of a basketball player are 6 feet above the ground, what is the angle of elevation of the player's line of sight when shooting a free throw to the rim of the basket?
-------------------------------------------------
Draw the picture.
You have a right triangle with base = 15 ft
and height = (10-6) = 4 ft
----
The angle of elevation is tan^-1(4/15) = 14.93 degrees.
==================
Cheers,
Stan H.